pg. 261
94) Swimming space. The length of a rectangular swimming pool is 2x - 1 meters, and the width is x + 2 meters. Write a polnomial that represents the area. Find the area if x is 5 meters.
answer: 2x^2+3x-2sq. m. 63 sq. m.

Answer by ilana(307)

I am guessing you are just checking, so yep, nice work!:)
On the off chance that you want to know how the book got that answer, for the area of a rectangle you just multiply length times width, so they multiplied using the FOIL method and then plugged in x=5.
I'll complete this one so you can see how it works.
We need to multiply length times width to get the area.
(2x-1)(x+2)
Use FOIL: First, Outside, Inside, Last
(2x)*(x) + (2x)*(2) + (-1)*(x) + (-1)*(2)
2x^2 + 4x + -1x + -2
2x^2 + 3x - 2
If x=5, this is 2*(5)^2 + 3*(5) - 2 = 2*25 + 15 - 2 = 65 - 2 = 63
Since this is area and we started with meters on both sides, we end up with square meters.